OpenStudy (anonymous):

The price of a baseball Ticket at Yankee Stadium increased by 2 1/4 over the last 3 years. If the original price of a ticket was $60, what is the new price of the ticket today?

OpenStudy (anonymous):

I answered 195, but I think it was probably 135....

Directrix (directrix):

> increased by 2 1/4 I am uncertain what this means. Could it be increased by 2.25 % ? Could it mean that the ticket price was multiplied by a factor of 2.25? -------------------

Directrix (directrix):

Let P be the original cost of the ticket. Let CP be the current cost of a ticket (3 years later) CP = P + 2.25*P CP = 3.25*P CP = 3.25 * 60 CP = $195.00

Directrix (directrix):

@ccmom82 I got the same answer you did.